OpenStudy (anonymous):

Given equations A and B as 2/5x+y=12 and 5/2y-x=6, respectively, which expression will eliminate the variable x?

OpenStudy (anonymous):

2/5x + y = 12 -x + 5/2y = 6 you would multiply the second equation by 2/5, and this will cancel out the x's

OpenStudy (anonymous):

what does it mean by " which expression " ? Because I know that if you multiply the second equation by 2/5, this will cause the x's to cancel out

OpenStudy (anonymous):

the second equation, when multiplied by 2/5. is : -2/5x + y = 12/5